# EMERGE: Integrating RAG for Improved Multimodal EHR Predictive Modeling

### v1

@@ Abstract
The integration of multimodal Electronic Health Records (EHR) data has notably advanced clinical predictive capabilities. However, current models that utilize clinical notes and multivariate time-series EHR data often lack the necessary medical context for precise clinical tasks. Previous methods using knowledge graphs (KGs) primarily focus on structured knowledge extraction. To address this, we propose EMERGE, a Retrieval-Augmented Generation (RAG) driven framework aimed at enhancing multimodal EHR predictive modeling. Our approach extracts entities from both time-series data and clinical notes by prompting Large Language Models (LLMs) and aligns them with professional PrimeKG to ensure consistency. Beyond triplet relationships, we include entities’ definitions and descriptions to provide richer semantics. The extracted knowledge is then used to generate task-relevant summaries of patients’ health statuses. These summaries are fused with other modalities utilizing an adaptive multimodal fusion network with cross-attention. Extensive experiments on the MIMIC-III and MIMIC-IV datasets for in-hospital mortality and 30-day readmission tasks demonstrate the superior performance of the EMERGE framework compared to baseline models. Comprehensive ablation studies and analyses underscore the efficacy of each designed module and the framework’s robustness to data sparsity. EMERGE significantly enhances the use of multimodal EHR data in healthcare, bridging the gap with nuanced medical contexts crucial for informed clinical predictions.
footnotetext: Equal contribution, Corresponding author.
@@
Keywords electronic health record, multimodal learning, large language model, retrieval-augmented generation
1 Introduction
@@
The advent of Electronic Health Records (EHR) marks a pivotal advancement in the way patient data is gathered and analyzed, contributing to a more effective and informed healthcare delivery system for clinical prediction [1, 2, 3, 4, 5]. This advancement is largely attributed to the utilization of multimodal EHR data, which primarily includes clinical notes and multivariate time-series data from patient records [6, 7, 8]. Such data types are integral to healthcare prediction tasks, mirroring the holistic approach practitioners adopt by leveraging various patient data points to inform their clinical decisions and treatment strategies, rather than depending on a single data source [9]. Deep learning-based methods have become the mainstream approach, processing multimodal data to learn a mapping from heterogeneous inputs to output labels [10, 11, 6]. However, in contrast to healthcare professionals, who have a deep understanding of medical contexts through extensive experience and knowledge, neural networks trained from scratch lack these insights into medical concepts [12]. Without deliberate integration of external knowledge, these networks often lack the ability or sensitivity to recognize crucial disease entities or laboratory test results within the EHR, essential for accurate prediction tasks [13]. In response, some recent studies have begun incorporating knowledge graphs to infuse additional medical insights into their analyses [14, 15]. These graphs offer a supplementary layer of clinically relevant concepts, thereby enhancing the model’s ability to provide contextually meaningful representations and interpretable evidence [16]. Despite these advancements, significant limitations remain in fully linking external knowledge with multiple EHR modalities, underscoring the imperative need for continuous research to integrate multi-source insights and improve the multimodal EHR data predictive modeling.
@@
Previous methods integrating external medical knowledge into EHR data analysis tend to extract knowledge from data modalities such as ICD disease codes, patient conditions, procedures, and drugs, neglecting the use of clinical notes and time-series data, which are more common and practical [17] (Limitation 1). Additionally, these methods primarily extract hierarchical and structured knowledge from clinical-context knowledge graphs. However, these medical concepts—entity names and their relationships into a graph have limited direct contribution to predictive tasks (Limitation 2). With Large Language Models (LLMs) like GPT-4 [18] demonstrating strong capabilities in diverse clinical tasks [13, 19, 20] and serving as large medical knowledge graphs (KGs) [21]. By prompting the LLM, GraphCare [22] constructs a GPT-KG using structured condition, procedure, and drug record data, represented as triples (entity 1, relationship, entity 2). It further employs graph neural networks for downstream tasks. However, this approach encounters the hallucination issue [23], where LLMs may generate incorrect or fabricated information. To mitigate this, GraphCare collaborates with medical professionals to scrutinize and remove potentially harmful content, a process that is both complex and labor-intensive, requiring significant expertise to validate and refine the generated triples. Moreover, directly generating the KG via LLMs introduces a domain gap since this task is likely untrained for the LLMs, leading to potentially lower accuracy compared to professional knowledge graphs built through established methodologies (Limitation 3).
@@
To overcome these limitations, we propose utilizing LLMs in a Retrieval-augmented Generation (RAG) approach [24]. The RAG framework integrates structured time-series EHR data, unstructured clinical notes, and an established KG (PrimeKG [25]) with LLM’s semantic reasoning capabilities [26]. The LLMs are prompted to generate comprehensive summaries of patients’ health statuses, and these summaries are then fused for downstream tasks. Despite its apparent simplicity, applying this method to clinical tasks presents several technical challenges:
@@
Challenge 1: How to extract entities from multimodal EHR data and match these entities with external KG consistently? Extracting entities from the diverse and complex formats of EHR data (including clinical notes and multivariate time-series data) is challenging. Moreover, unlike structured codes where it can directly compare the code-related entities’ embedding with KG’s entity, the entities extracted by LLM have hallucination issues. Accurately matching extracted entities with those in an external knowledge graph while eliminating the potential for hallucinations posed by LLMs is crucial for maintaining the integrity and reliability of the clinical prediction tasks [27].
@@
Challenge 2: How to encode and incorporate long-text retrieved knowledge with task-relevant characteristics? The extracted textual knowledge likely contains too many tokens [28] for conventional language model inputs (e.g., BERT supports only 512 tokens [29]). However, with the development of long-context LLMs [30], it is feasible to leverage LLMs to distill this knowledge further. Additionally, simply integrating the retrieved knowledge may not be task-specific, creating a gap between the knowledge and downstream tasks [31, 32, 33]. Therefore, a task-relevant prompting strategy [34] is necessary during the LLM distillation process.
@@
    Experimentally, our extensive experiments on the MIMIC-III and MIMIC-IV datasets, focusing on in-hospital mortality and 30-day readmission tasks, demonstrate EMERGE’s superior performance and the effectiveness of each designed module. Additionally, to meet practical clinical needs, we evaluate the model’s robustness with fewer training samples, showing EMERGE’s remarkable resilience against data sparsity. Moreover, a case study on the generated summaries reflects the framework’s soundness, also serving as an interpretable decision-making reference.
2 Related Work
2.1 Multimodal EHR Learning
@@
The evolution of medical technology has enabled the analysis of various medical modalities—ranging from clinical notes and time-series laboratory test data to demographics, conditions, procedures, drugs, and medical imaging. Noteworthy efforts in multimodal learning for healthcare include MedGTX [35] and M3Care [6]. MedGTX introduces a pre-trained model for joint multi-modal representation learning of structured and textual EHR data by interpreting structured EHR data as a graph and employing a graph-text multi-modal learning framework. M3Care compensates for the missing modalities by imputing task-related information in the latent space through auxiliary information from similar patients. M3Care leverages a task-guided modality-adaptive similarity metric to effectively handle missing modalities without relying on unstable generative models. The work of Zhang et al. [8] further explored the irregularity of time intervals in time-series EHR data and clinical notes via a time attention mechanism. Notably, Xu et al. [9] introduced a joint learning approach from visit sequences and clinical notes, employing Gromov-Wasserstein Distance for contrastive learning and dual-channel retrieval to enhance patient similarity analysis. Lee et al. [36] proposed a unified framework for learning across all EHR modalities, eschewing separate imputation modules in favor of modality-aware attention mechanisms.
Although the methods mentioned above perform well across multiple joint modalities, a common drawback is their limited consideration of incorporating clinical background information, wherein external medical knowledge could provide significant insights into the EHR data. Furthermore, the absence of semantic medical knowledge renders the training-from-scratch pipeline more challenging to converge, especially when data is scarce in practical clinical settings.
2.2 Incorporating External Knowledge for EHR
@@
Addressing the need to blend clinical background knowledge with EHR data, numerous studies have leveraged medical knowledge graphs (KGs) to enhance the EHR data representation learning process, thereby augmenting predictive performance. Techniques such as utilizing the ancestor information of nodes within KGs have been employed to refine medical representation learning, as seen in GRAM [10], which integrates hierarchical medical ontologies via a graph attention network. KAME [11] builds on this by embedding ontology information throughout the prediction process, enriching the contextual understanding of models. MedPath [14] employs graph neural networks to capture and integrate high-order connections from knowledge graphs into input representations, thereby enhancing the relevance and utility of external knowledge. MedRetriever [37] enhances health risk prediction and interpretability by leveraging unstructured medical text from authoritative sources. It combines EHR embeddings with features from target disease documents to retrieve relevant text segments. Collaborative graph learning models, such as CGL [38], explore patient-disease interactions and domain knowledge, while KerPrint [16] focus on addressing knowledge decay on multiple time visits. The advent of Large Language Models (LLMs) as comprehensive knowledge bases [21] offers new possibilities, exemplified by GraphCare [22], which creates a KG from structured EHR data for GNN learning, though it faces challenges related to content hallucination.
These efforts mostly concentrate on extracting knowledge from structured medical data, overlooking the rich semantic information embedded in unstructured EHR data. This oversight limits the potential for fully leveraging the depth of knowledge contained within EHRs, highlighting the need for methodologies that encompass both structured and unstructured data modalities.
3 Problem Formulation
3.1 EHR Datasets Formulation
@@
3.2 Predictive Objective Formulation
@@
$\hat{y}=\text{Framework}(\bm{x}_{TS},\bm{x}_{Note},KG)$ (1)
@@
3.3 Notation Table
The notations and their descriptions used in this paper are presented in Table 1.
Table 1: Notations symbols and their descriptions
Notations Descriptions
$N$ Number of patients
$KG$ External knowledge graphs
$\bm{x}_{TS}$ Time-series data of one patient
$\bm{x}_{Note}$ Clinical note of one patient
$\bm{x}_{RAG}$ Retrieved textual knowledge of one patient
$T$ Number of visits for a patient in time-series EHR data
$F$ Number of features in time-series EHR data
$\bm{h}_{\square}$ Representation of modality or fused hidden states
$E_{TS}$ Entity set extracted from a single time-series EHR data
$E_{Note}$ Entity set extracted from one clinical note
$\theta$ Cosine similarity between two embedding vectors
$\epsilon$ Threshold for identifying anomalies in time-series data
$\eta$ Threshold for matching extracted entities with nodes in knowledge graph
$s_{i,t}$ Z-score value for the $i$-th feature at visit $t$ of one patient
$\boldsymbol{W}_{\square}$ Parameter matrices of linear layers. Footnote $\square$ denotes the name of the layer
$\bm{z}_{i}$ Fused final representation of the $i$-th patient
LM Language Model (basically BERT-based model)
LLM Large Language Model (basically GPT-based model)
4 Methodology
Figure 1: Overall architecture of our proposed EMERGE framework. The modules enclosed within the dashed box illustrate the RAG-driven enhancement pipeline.
Figure 1 shows the overall framework architecture of EMERGE. It consists of three main modules:
  • Multimodal Embedding Extraction applies GRU as time-series data’s encoder and pretrained long-context language model for clinical notes.
  • RAG-Driven Enhancement Pipeline retrieves entities and links knowledge in established KG for both time-series and clinical notes modalities. We integrate entities triples relationships along with their definitions and descriptions along with disease relationships into a prompt. The prompt serves as the instruction for the LLM to generate the patient health status summary.
  • Multimodal Fusion Network incorporates the generated summaries and adaptively fuses multiple modalities for further downstream tasks.
4.1 Multimodal EHR Embedding Extraction
@@
When dealing with time-series data, we employ the Gated Recurrent Unit (GRU) network as the encoder. GRU is a highly efficient variant of recurrent neural networks, capable of capturing the time dependencies in sequence data and encoding this temporally linked information. We extract the representation of time-series as below:
$\bm{h}_{TS}={\rm GRU}(x_{TS}).$ (2)
where $x_{TS}$ is the time-series data and $h_{TS}$ denotes the output of time-series encoder.
As for text records, we utilize medical domain language model to obtain text embeddings, represented as TextEncoder. Formally,
$\bm{h}_{Note}={\rm TextEncoder}(\bm{x}_{Note}).$ (3)
where $\bm{x}_{Note}$ is the textual clinical notes and ${h}_{Note}$ denotes the note representation.
4.2 RAG-Driven Enhancement Pipeline
4.2.1 Extract Entities from Multimodal EHR Data
In order to exploit the expert information encapsulated within the knowledge graph, it is necessary to extract disease entities from both time-series data and clinical notes, and subsequently align them with the information present in the graph. The set of disease entities in the time-series data is denoted as $\bm{E}_{TS}$, while those in the clinical notes data are denoted as $\bm{E}_{Note}$. Naturally, we design two separate processes tailored to each modality.
@@ 4.2.1 Extract Entities from Multimodal EHR Data
Time-series data is a structured format encompassing feature names and resultant values post clinical examination. Each feature name reflects specific aspects of an individual’s physical condition, highlighting the deviations from the reference range. As shown in Figure 2, the specified record showcases a low blood pressure and high blood urea nitrogen, significantly surpassing the normal range. This imply the potential risk of hypo-tension and uremia for the patient. Indeed, such feature names occur in diseases definitions and descriptions, typically indicating serious health threats.
For each patient, there are usually more than one entity (or abnormal feature), and some may be missing values. Consequently, our focus is primarily on non-empty values. For each feature $\bm{x}_{TS_{i}}$, we can identify outliers through the z-score method  [39], which measures anomalies by calculating the deviation of data points from the mean, using standard deviation as a unit as below:
${s_{i}}=\frac{\bm{x}_{TS_{i}}-{\rm mean}(\bm{x}_{TS_{i}})}{{\rm std}(\bm{x}_{TS_{i}})}$ (4)
@@
Figure 2: Process of information retrieval for time-series data.
@@
Contrary to structured data, clinical notes are presented in a textual format, which makes it challenging to comprehend and extract valuable information. However, LLMs have exhibited exceptional performance on natural language understanding tasks, including named entity recognition (NER). Therefore, we utilize an LLM to identify potential disease names that the patient may have, as shown in Figure 4. Moreover, we implement specified rules for effective post-processing.
  1. 1.
    Entities Extraction: We employ an example and clear instruction in prompt, instructing LLM to concentrate on entities of disease that the patient may suffer from. The complete prompt template is in Figure 3. Sometimes there may be no entities yield in single invocation, so we utilize multiple rounds to incrementally expand the current extracted entity set as below:
    $\bm{E}^{i}_{Note}$ $={\rm LLM}(concat(\bm{P}_{Extract},\bm{x}_{Note}))$ (5)
    $\bm{E}_{Note}$ $\leftarrow\bm{E}_{Note}\bigcup\bm{E}^{i}_{Note}$ (6)
@@
    Figure 3: Prompt template for extracting entities.
  2. 2.
    Entities Refinement: Considering the hallucination issue associated with LLM, we design a post-processing process to address it. This process consists of three primary steps: firstly, we discard entities that do not appear in the original text; secondly, we leverage LLM to filter entities not in disease type; and finally, we delete duplicated entities to prevent semantic redundancy.
    $\bm{E}_{Note}\leftarrow\bm{E}_{Note}-\bm{E}_{illegal}$ (7)
@@
To ensure the quantity and quality of the extracted entities, we execute step 1 and step 2 iteratively until achieving convergence.
Figure 4: Process of information retrieval for textual clinical notes. The grey block in potential diseases means no corresponding node found in external KG.
4.2.2 Retrieve Information from External KG

### v2

# EMERGE: Enhancing Multimodal Electronic Health Records Predictive Modeling with Retrieval-Augmented Generation

@@ Abstract.
The integration of multimodal Electronic Health Records (EHR) data has significantly advanced clinical predictive capabilities. Existing models, which utilize clinical notes and multivariate time-series EHR data, often fall short of incorporating the necessary medical context for accurate clinical tasks, while previous approaches with knowledge graphs (KGs) primarily focus on structured knowledge extraction. In response, we propose EMERGE, a Retrieval-Augmented Generation (RAG) driven framework to enhance multimodal EHR predictive modeling. We extract entities from both time-series data and clinical notes by prompting Large Language Models (LLMs) and align them with professional PrimeKG, ensuring consistency. In addition to triplet relationships, we incorporate entities’ definitions and descriptions for richer semantics. The extracted knowledge is then used to generate task-relevant summaries of patients’ health statuses. Finally, we fuse the summary with other modalities using an adaptive multimodal fusion network with cross-attention. Extensive experiments on the MIMIC-III and MIMIC-IV datasets’ in-hospital mortality and 30-day readmission tasks demonstrate the superior performance of the EMERGE framework over baseline models. Comprehensive ablation studies and analysis highlight the efficacy of each designed module and robustness to data sparsity. EMERGE contributes to refining the utilization of multimodal EHR data in healthcare, bridging the gap with nuanced medical contexts essential for informed clinical predictions. We have publicly released the code at https://github.com/yhzhu99/EMERGE.
@@
electronic health record; multimodal learning; large language model; retrieval-augmented generation
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679582ccs: Applied computing Health informaticsccs: Information systems Data mining
1. Introduction
@@
The advent of Electronic Health Records (EHR) marks a pivotal advancement in the way patient data is gathered and analyzed, contributing to a more effective and informed healthcare delivery system for clinical prediction (Gao et al., 2024; Ma et al., 2023; Liao et al., 2024). This advancement is largely attributed to the utilization of multimodal EHR data, which primarily includes clinical notes and multivariate time-series data from patient records (Zhang et al., 2022; Wang et al., 2024; Zhang et al., 2023a). Such data types are integral to healthcare prediction tasks, mirroring the holistic approach practitioners adopt by leveraging various patient data points to inform their clinical decisions and treatment strategies, rather than depending on a single data source (Xu et al., 2023). Deep learning-based methods have become the mainstream approach, processing multimodal data to learn a mapping from heterogeneous inputs to output labels (Choi et al., 2017; Ma et al., 2018; Zhang et al., 2022). However, in contrast to healthcare professionals, who have a deep understanding of medical contexts through extensive experience and knowledge, neural networks trained from scratch lack these insights into medical concepts (Miotto et al., 2018). Without deliberate integration of external knowledge, these networks often lack the ability or sensitivity to recognize crucial disease entities or laboratory test results within the EHR, essential for accurate prediction tasks (Zhu et al., 2024a). In response, some recent studies have begun incorporating knowledge graphs to infuse additional medical insights into their analyses (Ye et al., 2021a; Gao et al., 2022). These graphs offer a supplementary layer of clinically relevant concepts, thereby enhancing the model’s ability to provide contextually meaningful representations and interpretable evidence (Yang et al., 2023). Despite these advancements, significant limitations remain in fully linking external knowledge with multiple EHR modalities, underscoring the imperative need for continuous research to integrate multi-source insights and improve the multimodal EHR data predictive modeling.
@@
Previous methods integrating external medical knowledge into EHR data analysis tend to extract knowledge from data modalities such as ICD disease codes, patient conditions, procedures, and drugs, neglecting the use of clinical notes and time-series data, which are more common and practical (Rajkomar et al., 2018) (Limitation 1). Additionally, these methods primarily extract hierarchical and structured knowledge from clinical-context knowledge graphs. However, these medical concepts—entity names and their relationships into a graph have limited direct contribution to predictive tasks (Limitation 2). With Large Language Models (LLMs) like GPT-4 (Achiam et al., 2023) demonstrating strong capabilities in diverse clinical tasks (Zhu et al., 2024a; Wornow et al., 2023; Shi et al., 2024) and serving as large medical knowledge graphs (KGs) (Sun et al., 2023). By prompting the LLM, GraphCare (Jiang et al., 2024) constructs a GPT-KG using structured condition, procedure, and drug record data, represented as triples (entity 1, relationship, entity 2). It further employs graph neural networks for downstream tasks. However, this approach encounters the hallucination issue (Zhang et al., 2023b), where LLMs may generate incorrect or fabricated information. To mitigate this, GraphCare collaborates with medical professionals to scrutinize and remove potentially harmful content, a process that is both complex and labor-intensive, requiring significant expertise to validate and refine the generated triples. Moreover, directly generating the KG via LLMs introduces a domain gap since this task is likely untrained for the LLMs, leading to potentially lower accuracy compared to professional knowledge graphs built through established methodologies (Limitation 3).
@@
To overcome these limitations, we propose utilizing LLMs in a Retrieval-augmented Generation (RAG) approach (Lewis et al., 2020). The RAG framework integrates structured time-series EHR data, unstructured clinical notes, and an established KG (PrimeKG (Chandak et al., 2023)) with LLM’s semantic reasoning capabilities (Wang et al., 2023). The LLMs are prompted to generate comprehensive summaries of patients’ health statuses, and these summaries are then fused for downstream tasks. Despite its apparent simplicity, applying this method to clinical tasks presents several technical challenges:
@@
Challenge 1: How to extract entities from multimodal EHR data and match these entities with external KG consistently? Extracting entities from the diverse and complex formats of EHR data (including clinical notes and multivariate time-series data) is challenging. Moreover, unlike structured codes where it can directly compare the code-related entities’ embedding with KG’s entity, the entities extracted by LLM have hallucination issues. Accurately matching extracted entities with those in an external knowledge graph while eliminating the potential for hallucinations posed by LLMs is crucial for maintaining the integrity and reliability of the clinical prediction tasks (Imrie et al., 2023).
@@
Challenge 2: How to encode and incorporate long-text retrieved knowledge with task-relevant characteristics? The extracted textual knowledge likely contains too many tokens (Xiao et al., 2023) for conventional language model inputs (e.g., BERT supports only 512 tokens (Devlin et al., 2018)). However, with the development of long-context LLMs (Zhu et al., 2024c), it is feasible to leverage LLMs to distill this knowledge further. Additionally, simply integrating the retrieved knowledge may not be task-specific, creating a gap between the knowledge and downstream tasks (Bai et al., 2023b, 2024, c). Therefore, a task-relevant prompting strategy (OpenAI, 2023) is necessary during the LLM distillation process.
@@
    Experimentally, our extensive experiments on the MIMIC-III and MIMIC-IV datasets, focusing on in-hospital mortality and 30-day readmission tasks, demonstrate EMERGE’s superior performance and the effectiveness of each designed module. Additionally, to meet practical clinical needs, we evaluate the model’s robustness with fewer training samples, showing EMERGE’s remarkable resilience against data sparsity.
2. Related Work
2.1. Multimodal EHR Learning
@@
Advances in medical technology enable analysis of various medical modalities, including clinical notes, time-series lab data, demographics, conditions, procedures, drugs, and imaging. MedGTX (Park et al., 2022) introduces a pre-trained model for joint multi-modal representation learning, interpreting structured data as a graph and using a graph-text multi-modal framework. M3Care (Zhang et al., 2022) addresses missing modalities by imputing task-related information in the latent space with auxiliary data from similar patients, employing a modality-adaptive similarity metric to handle missing data.  Zhang et al. (2023a) explore irregular time intervals in time-series EHR data and clinical notes via a time attention mechanism.  Xu et al. (2023) propose a joint learning approach from visit sequences and clinical notes, using Gromov-Wasserstein Distance for contrastive learning and dual-channel retrieval to enhance patient similarity analysis.  Lee et al. (2023) introduce a unified framework for learning across all EHR modalities with modality-aware attention mechanisms, avoiding separate imputation modules.
Despite their effectiveness, these methods often overlook clinical background information, where external medical knowledge could enhance EHR data insights. The absence of semantic medical knowledge also complicates the training pipeline, especially with limited data.
2.2. Incorporating External Knowledge for EHR
@@
To integrate clinical knowledge with EHR data, several studies leverage medical knowledge graphs (KGs) to enhance EHR representation learning and predictive performance. GRAM (Choi et al., 2017) uses hierarchical medical ontologies via a graph attention network to refine medical representations. KAME (Ma et al., 2018) embeds ontology information throughout the prediction process, enriching contextual understanding. MedPath (Ye et al., 2021a) employs graph neural networks to integrate high-order connections from KGs into input representations. MedRetriever (Ye et al., 2021b) enhances health risk prediction and interpretability by combining EHR embeddings with features from disease-specific documents. Collaborative graph learning models like CGL (Lu et al., 2021) explore patient-disease interactions and domain knowledge, while KerPrint (Yang et al., 2023) addresses knowledge decay across multiple visits. Recent advancements in Large Language Models (LLMs) as comprehensive knowledge bases (Sun et al., 2023) offer new possibilities, as seen in GraphCare (Jiang et al., 2024), which creates a KG from structured EHR data for GNN learning, despite challenges like hallucination.
These studies primarily focus on structured medical data, often neglecting the rich semantic information in unstructured EHR data. This limitation underscores the need for methods that comprehensively utilize both structured and unstructured data.
3. Problem Formulation
3.1. EHR Datasets Formulation
@@
3.2. Predictive Objective Formulation
@@
(1) $\hat{y}=\text{Framework}(\bm{x}_{TS},\bm{x}_{Note},KG)$
@@
4. Methodology
Figure 1. Overall architecture of our proposed EMERGE framework. The modules enclosed within the dashed box illustrate the RAG-driven enhancement pipeline. “LM” denotes Language Model (basically BERT-based model), while “LLM” in this paper normally refers to the GPT-based Large Language Model.
Figure 1 shows the overall framework architecture of EMERGE.
4.1. Multimodal EHR Embedding Extraction
@@
When dealing with time-series data, we employ the Gated Recurrent Unit (GRU) network as the encoder. GRU is a highly efficient variant of recurrent neural networks, capable of capturing the time dependencies in sequence data and encoding this temporally linked information. We extract the representation of time-series data as follows:
(2) $\bm{h}_{TS}={\rm GRU}(x_{TS}),$
where $x_{TS}$ is the time-series data and $h_{TS}$ denotes the output of the time-series encoder.
As for text records, we utilize a medical domain language model to obtain text embeddings, represented as TextEncoder:
(3) $\bm{h}_{Note}={\rm TextEncoder}(\bm{x}_{Note}).$
where $\bm{x}_{Note}$ are the textual clinical notes and ${h}_{Note}$ denotes the note representation.
4.2. RAG-Driven Enhancement Pipeline
4.2.1. Extract Entities from Multimodal EHR Data
To exploit the expert information encapsulated within the knowledge graph, it is necessary to extract disease entities from both time-series data and clinical notes, and subsequently align them with the information present in the graph. The set of disease entities in the time-series data is denoted as $\bm{E}_{TS}$, while those in the clinical notes data are denoted as $\bm{E}_{Note}$. Naturally, we design two separate processes tailored to each modality.
@@ 4.2.1. Extract Entities from Multimodal EHR Data
Time-series data is a structured format encompassing feature names and resultant values post-clinical examination. Each feature name reflects specific aspects of an individual’s physical condition, highlighting the deviations from the reference range. As shown in Figure 2, the specified record showcases low blood pressure and high blood urea nitrogen, significantly surpassing the normal range. This implies the potential risk of hypotension and uremia for the patient. Indeed, such feature names occur in disease definitions and descriptions, typically indicating serious health threats.
For each patient, there are usually more than one entity (or abnormal feature), and some may be missing values. Consequently, our focus is primarily on non-empty values. For each feature $\bm{x}_{TS_{i}}$, we can identify outliers through the z-score method (Curtis et al., 2016), which measures anomalies by calculating the deviation of data points from the mean, using standard deviation as a unit as below:
(4) $s_{i}=\frac{\bm{x}_{TS_{i}}-{\rm mean}(\bm{x}_{TS_{i}})}{{\rm std}(\bm{x}_{TS_{i}})}$
@@
Figure 2. Process of information retrieval for time-series data.
@@
Unlike structured data, clinical notes are presented in a textual format, which makes it challenging to comprehend and extract valuable information. However, LLMs have exhibited exceptional performance on natural language understanding tasks, including named entity recognition (NER). Therefore, we utilize an LLM to identify potential disease names that the patient may have, as shown in Figure 4. Moreover, we implement specified rules for effective post-processing.
Figure 3. Prompt template for extracting entities.
  1. (1)
    Entities Extraction: We include an example and provide clear instructions in the prompt (Figure 3), instructing the LLM to concentrate on disease entities that the patient may suffer from. Sometimes, there may be no entities yielded in a single invocation, so we utilize multiple rounds to incrementally expand the current extracted entity set as shown below:
    (5) $\bm{E}^{i}_{Note}$ $={\rm LLM}(concat(\bm{P}_{Extract},\bm{x}_{Note}))$
    (6) $\bm{E}_{Note}$ $\leftarrow\bm{E}_{Note}\bigcup\bm{E}^{i}_{Note}$
@@
  2. (2)
    Entities Refinement: Considering the hallucination issue associated with LLMs, we design a post-processing process to address it. This process consists of three primary steps: first, we discard entities that do not appear in the original text; second, we leverage an LLM to filter entities not in the disease type; and finally, we delete duplicated entities to prevent semantic redundancy.
    (7) $\bm{E}_{Note}\leftarrow\bm{E}_{Note}-\bm{E}_{illegal},$
@@
To ensure the quantity and quality of the extracted entities, we execute steps 1 and 2 iteratively until achieving convergence.
Figure 4. Process of information retrieval for textual clinical notes. The grey block in potential diseases means no corresponding node found in external KG.
4.2.2. Retrieve Information from External KG